### v3

# fold/.style=folder, grow’=0, s sep=0.1pt, arr/.style=edge=-latex

Abstract
This paper presents a deep reinforcement learning approach for synthesizing unitaries into quantum circuits. Unitary synthesis aims to identify a quantum circuit that represents a given unitary while minimizing circuit depth, total gate count, a specific gate count, or a combination of these factors. While past research has focused predominantly on continuous gate sets, synthesizing unitaries from the parameter-free Clifford+T gate set remains a challenge. Although the time complexity of this task will inevitably remain exponential in the number of qubits for general unitaries, reducing the runtime for simple problem instances still poses a significant challenge. In this study, we apply the tree-search method Gumbel AlphaZero to solve the problem for a subset of exactly synthesizable Clifford+T unitaries. Our approach can synthesize unitaries for up to five qubits generated from the set of randomized quantum circuits with up to $60$ gates. Furthermore, our inference times are around $30$ seconds on a single GPU on average, surpassing state-of-the-art algorithms QuantumCircuitOpt and MIN-T-SYNTH for higher qubit numbers. Our work provides a competitive baseline for synthesis algorithms to be developed in the upcoming years.
Index Terms:
Unitary synthesis, Reinforcement learning, MCTS, AlphaZero
I Introduction
The advent of quantum computing foreshadows a paradigm shift in our computational capabilities, sparking significant efforts towards scalable implementations of quantum algorithms on actual quantum hardware. Like classical computers, executing a quantum algorithm on a device requires converting a high-level algorithmic representation into low-level, hardware-dependent operations that can be executed on the device. The umbrella term for this task is quantum circuit compilation, a broad set of different constituents, encompassing optimization, qubit mapping and routing, gate scheduling, error correction, and classical control integration, among others [1].
One of the sub-tasks of quantum compilation for a problem-specific unitary matrix is to identify a quantum circuit comprised of gates from a pre-defined gate set whose corresponding unitary matrix closely approximates or precisely matches it, known as approximate and exact unitary synthesis respectively. Our focus in this paper is on synthesizing quantum circuits from unitary matrices, an important sub-routine in any quantum development kit, that can generate efficient quantum circuits for targeted hardware [2, 3, 4].
In the fault-tolerant setting [5], the Clifford+T gate set plays a crucial role due to its efficiency in realizing fault-tolerant quantum computation. The T-gate is essential for enabling universal quantum computing, but it is considered costly because of the necessity of magic state distillation [6], which introduces significant resource overhead and increases circuit depth [7]. Similarly, in noisy intermediate-scale quantums (NISQs) devices [8], two-qubit entangling gates are typically more error-prone and have longer execution times compared to single-qubit gates [9]. To reduce overall error rates and improve the reliability of synthesized circuits, synthesis algorithms must take into account the specific characteristics of certain gates, such as the T-gate in fault-tolerant circuits and two-qubit gates in NISQ devices. By minimizing the use of these costly gates, the synthesis algorithms can generate circuits that are more resilient to errors and better suited for their respective computing paradigms.
Although exact and provably optimal synthesis methods for Clifford+T circuits exist [10, 11], their practicality is limited due to runtime complexities that grow exponentially with the number of qubits and the number of T-gates. This limitation highlights the need to explore heuristic approaches to make unitary synthesis more applicable to real-world problems where the circuit representation is unknown, such as problem-specific quantum oracles. Recent developments in reinforcement learning (RL) have shown that it is a highly successful framework for learning such heuristics through directed trial and error. Inspired by initial contributions towards RL-based synthesizers [12], we model unitary synthesis as a sequential decision-making problem in which an RL agent appends one gate at a time to the quantum circuit. Although model-free RL has proven successful on simple tasks [12, 13], its application remains limited to circuits with a small number of qubits and gate sets.
In recent years, model-based RL utilizing Monte Carlo tree search (MCTS) [14, 15], such as AlphaGo [16] and AlphaZero [17], have shown great promise in solving complex problems with large search spaces [18, 19, 20]. Recently, the Gumbel AlphaZero algorithm [21] has proven effective in solving complex problems while requiring fewer computational resources compared to its predecessors.
To address the limitations of past unitary synthesis methods, we propose a novel method that leverages the strengths of Gumbel AlphaZero and incorporates domain-specific knowledge to enable efficient and scalable unitary synthesis. In our study, we address the exact synthesis of Clifford+T unitaries, even though, in theory, our method is not limited to one specific gate set or an exact synthesis regime. We show that our method scales up to 5-qubit circuits and can, on average, synthesize unitaries into circuits comprised of fewer T-gates than those present in the circuits used to generate the target unitary. Furthermore, when the agent is tasked with synthesizing structured unitaries such as the Toffoli and Peres gates, we show that it performs optimally in terms of the number of T-gates for almost all examples.
In contrast to other approaches that are trained for one specific unitary [13, 22], our agent can synthesize a wide range of input unitaries without the need for retraining, highlight the effectiveness and generalizability of our Gumbel AlphaZero-based approach. To the best of the authors’ knowledge, we are the first to scale RL-based unitary synthesis past the one- or two-qubit setting with a generalizing agent.
The paper is organized as follows: Section II introduces unitary synthesis and RL. Section III reviews previous work in unitary synthesis using classical heuristics and RL methods. Section V describes our RL environment formulation. Section IV presents our Gumbel AlphaZero-based algorithm. Section VI details the experimental setup and discusses the results, and Section VII explores future improvements. The Appendix provides a brief overview of quantum circuit optimisation.
II Fundamentals
II-A Quantum Unitary Synthesis
Unitary synthesis is the process of devising an efficient quantum circuit that implements a given unitary matrix, which can then be processed by a quantum compiler for execution on a real device. Formally, for a gate set $\mathcal{G}$, a circuit $C$ is said to be over $\mathcal{G}$ if it consists solely of gates from $\mathcal{G}$. We denote the set of all such circuits as $\langle\mathcal{G}\rangle$. Given a unitary $U\in\mathcal{U}(2^{n})$, a gate set $\mathcal{G}$, and a precision $\epsilon$, the goal of unitary synthesis is to find a circuit $C\in\langle\mathcal{G}\rangle$ whose unitary representation $V_{C}$ satisfies $d(U,V_{C})\leq\epsilon$, where $d:\mathcal{U}(2^{n})\times\mathcal{U}(2^{n})\rightarrow\mathbb{R}$ is a distance metric. Setting $\epsilon$ to zero defines the problem of exact synthesis.
a) Distance metrics: The Hilbert-Schmidt distance, average gate fidelity, and trace distance are among the most commonly used metrics for comparing unitaries. A detailed discussion of these and other metrics used in quantum information theory is available in Ref. [23].
In our approach, we utilize the Hilbert-Schmidt inner product (HS), defined as $\langle U,V_{t}\rangle=\mathrm{Tr}(V_{t}^{\dagger}U)$, to measure the similarity between a target unitary $U$ and another unitary $V_{t}$, which represents the synthesized circuit at time step $t$. The cost function for optimization in a $2^{n}$-dimensional space, corresponding to $n$ qubits, is given by:
$C_{HS}=1-\frac{1}{4^{n}}|\text{Tr}(V_{t}^{\dagger}U)|^{2}$ (1)
Note that this cost function is invariant to the global phase, meaning unitaries are considered equivalent under this metric, even if they differ by a global phase.
b) Clifford+T: The Clifford+T gate set, essential for many fault-tolerant quantum computing protocols, includes the Clifford gates, which map Pauli strings to Pauli strings up to a phase of {±$\{\pm 1,\pm i\}$, and the T-gate. However, as already mentioned, implementing error-correcting codes in fault-tolerant quantum computers poses a challenge due to the T-gate, which necessitates minimizing its occurrences in a circuit [7]. In cases where an arbitrary unitary cannot be exactly synthesized with Clifford+T, a sequence of Clifford+T gates can still be found to approximate the unitary to any desired level of accuracy [24].
To quantify the optimality of T-gate usage, we adopt the T-count metric, $\mathcal{T}(U)$. It is defined as the minimum number of T-gates required to implement a unitary $U$ up to a global phase for unitaries that can be exactly synthesized from the Clifford+T gate set [25, 26, 27].
II-B Reinforcement Learning
Recent advancements in RL have led to its application in quantum computing for tasks such as quantum compilation, circuit optimization, and routing [12, 28, 29, 30, 31, 32, 33, 34, 35, 36].
RL is a learning paradigm, where an agent learns to optimize a reward signal through iteratively interacting with an environment along some time horizon. The agent can perform actions that may be discrete, continuous, or a combination of both. The problem is formally modelled as a Markov Decision Process (MDP), characterized by a set of states $\mathcal{S}$, a set of actions $\mathcal{A}$, a reward function $r(s,a,s^{\prime})$, a discount factor $\gamma$, and an environment dynamics function $p(s^{\prime}|s,a)$, describing the probability of transitioning to state $s^{\prime}$ given the current state $s$ and action $a$.
The primary goal is to devise an optimal policy $\pi^{*}(a|s)$, a probabilistic function mapping states to action probabilities, that maximizes the expected sum of discounted rewards, called the expected return, for every possible environment state. Given the reward trajectory of an episode, the return at time step $t$ is defined as $G_{t}=\sum_{k=t+1}^{T}\gamma^{k-t-1}r_{k}$, where $r_{k}$ is the reward at time $k$, and $T$ is the episode’s length. Given a policy $\pi$ and a general state $s$, we denote the expected return of $s$ under $\pi$ as $v_{\pi}(s)$, also known as the value function. The expected return of taking an action $a$ in a state $s$ under $\pi$, denoted $q_{\pi}(s,a)$, is referred to as the state-action or Q-value function. In the following, we work with estimators of Q-value function and denote it as ${Q}(s,a)$. For more details on RL, refer to Ref. [37].
III Related work
III-A Classical Unitary Synthesis
Pioneering theoretical contributions to the exact synthesis of single-qubit unitaries using the Clifford+T gate set were made by Kliuchnikov et al. [38]. Their work centered on determining the possibility of exact unitary synthesis and devising an efficient sequence of gates for implementation. Giles and Selinger further extended these concepts to general $n$-qubit unitaries, introducing an exact synthesis algorithm [26], which was later refined by Kliuchnikov [39]. Although these methods provided foundational insights, they did not prioritize T-gate reduction or gate count minimization, and both exhibited exponential run-time complexities. Selinger first proposed a highly efficient algorithm tailored for approximating arbitrary single-qubit unitaries with the Clifford+T gate set [40], marking a significant stride towards practical unitary synthesis.
The first deterministic method aimed at finding the T-count of a unitary which is exactly synthesizable using Clifford+T gate set was introduced by Gosset et al. [25]. This method leveraged the channel representation of unitaries and a decision algorithm for determining the T-count of a unitary $U$, denoted as $\mathcal{T}(U)\leq m$, for a specified integer $m$. The complexity of this algorithm was characterized as $\mathcal{O}(2^{nm}\text{poly}(m,2^{n}))$, with $n$ representing the number of qubits and $m$ the T-count.
The first heuristic algoroithm for synthesizing a T-count optimal unitary was developed by Mosca and Mukhopadhyay [11]. Their algorithm, MIN-T-SYNTH, prunes the search space by exploiting the structure of the channel representation of the unitary, achieving a conjectured polynomial complexity relative to the T-count. Despite its advancements, the algorithm’s exponential complexity relative to the number of qubits remains a challenge.
QuantumCircuitOpt [41] offers a practical approach to unitary synthesis by converting the circuit design task into specialized mixed-integer programs (MIP) [42], solvable using MIP solvers such as Gurobi [43]. This framework is versatile, accommodating various gate sets beyond Clifford+T, including parameterized gates and higher-level gates like controlled-V and Toffoli. However, exact synthesis with minimal T-gates is practically feasible only for low-depth circuits with a small number of qubits.
Finally, the work in Ref. [44] used decision diagram representation of a unitary matrix [45] to synthesize exactly representable unitaries with the Clifford+T gate set. Their algorithm was able to synthesize unitaries with up to 15 qubits and millions of T-gates.
III-B Reinforcement Learning-Based Unitary Synthesis
Given the large design space of quantum circuits, the automation of unitary synthesis via reinforcement learning has been the subject of various studies. An initial theoretical investigation into the feasibility of modelling quantum state preparation and gate synthesis as a Markov Decision Process (MDP) was conducted in Ref. [46], showing that optimal gate sequences could be discovered using this framework for both noisy and noiseless single-qubit scenarios.
RL has since been applied to a range of quantum circuit synthesis problems, particularly those involving parameterized quantum circuits. For example, double Q-learning was utilized for the variational compilation of specific 2 and 3-qubit unitaries in Ref. [13], including the controlled-H, quantum Fourier transform, Toffoli, and W gates. Additionally, Monte Carlo tree search methods (MCTS) have been employed for hybrid optimization and problem-specific quantum circuit design, as seen in Refs. [47, 48]. Other research has focused on hardware-specific quantum circuit design using reinforcement learning, as detailed in Refs. [49, 50, 51, 52].
A recent study has further demonstrated the use of RL in the synthesis of quantum circuits, specifically targeting state preparation [53]. This research explored the relationship between the depth of synthesized circuits, the depth of circuits used for initialization during training, and the number of qubits. The study showed promising results in synthesizing simple 2-qubit product and Bell states.
For approximate synthesis, Moro et al. pioneered the concept of synthesizing quantum circuits from fixed unitaries as a sequential decision-making process [12]. Employing DQN [54] and PPO [55], they successfully synthesized general single-qubit Haar-unitaries under the HRC efficient universal set [56] and a rotational gate set. Chen et al. extended this approach to two-qubit unitaries using a combination of DQN and A* search, albeit with a search complexity that grows exponentially with the number of gates required [57].
In this paper, we explore the application of reinforcement learning to unitary synthesis, aiming to extend its scalability to circuits with more qubits and greater depth. We benchmark our method against MIN-T-SYNTH [11] and QuantumCircuitOpt [41] using structured unitaries such as controlled-Hadamard, Toffoli, and Peres gates.
IV Gumbel AlphaZero
Although the AlphaZero algorithm [17] has demonstrated significant success, it can encounter difficulties when the MCTS simulation budget is limited, which may lead to under-representation of certain actions and an inadequate enhancement of the policy network that directs the search. To overcome this, the Gumbel AlphaZero framework [21] was developed, which excels in scenarios where the number of possible discrete actions $n_{a}$ in a certain state greatly exceeds the simulation budget $n_{sim}$.
Gumbel AlphaZero employs the Gumbel-Top-$k$ trick at the root node to sample a subset of $k$ actions from a total of $n_{a}$ possible actions without replacement. This method utilizes a vector $g\in\mathbb{R}^{n_{a}}$ of independent Gumbel-distributed random variables and a policy network $\pi_{\theta}$, where $\pi_{\theta}(s_{0})=\text{logits}\in\mathbb{R}^{n_{a}}$ for the root state $s_{0}$. The Gumbel-Top-$k$ trick selects the actions that maximize the expression $g+\text{logits}$, denoted as $\mathrm{argtop}\;(g+\text{logits},k)$, where $\mathrm{argtop}$ returns the indices of the $k$ largest values of the vector. Actions are then chosen from the sub-sampled set according to
[…]
where $Q(s_{0},a)\in\mathbb{R}^{n_{a}}$ represents the empirical state-action values, $\sigma$ is any arbitrary monotonically increasing function, and $A$ is the action selected. The authors prove that selecting $A_{n_{sim}+1}$ as per Eq. 2 guarantees policy improvement, given the estimate of the state-action values for the root state is accurate. They instantiate $\sigma$ as
$\sigma(Q(s_{0},a))=(c_{\text{visit}}+\underset{a^{\prime}}{\mathrm{max}}N(s_{0},a^{\prime}))c_{\text{scale}}Q(s_{0},a)$ (3)
inspired by the MPO policy update [58], where $N(s_{0},a)$ denotes the visitation count for action $a$ at root state $s$, and $c_{\text{visit}}$ and $c_{\text{scale}}$ are hyper-parameters. Additionally, sequential halving [59] is applied to efficiently allocate the simulation budget across the sampled actions, offering a bandit algorithm that focuses on minimizing simple regret, in contrast to the cumulative regret minimization in AlphaZero. To reduce variance in the Q-value estimates, action selection at non-root nodes is deterministic, diverging from the stochastic selection used in the standard AlphaZero formulation.
V Our RL framework
Since we focus on exact synthesis, we assume from the outset that the target unitary can be exactly synthesized using the Clifford+T gate set. Our corresponding RL framework is depicted in Fig. 1 and, in its essence, aligns with the model proposed by Moro et al. [12].
a) MDP Definition: The agent is tasked with the iterative construction of a quantum circuit that embodies a unitary $V$, which should coincide with a predefined target unitary $U$. In the following, we denote by $V_{t}$ the unitary representation of the current circuit $C_{t}$ at time step $t$. Actions within our framework correspond to the selection and addition of a gate from the gate set to the circuit at the earliest vacant moment for the required qubit(s). Each potential placement of a gate constitutes a distinct action. For instance, with single-qubit gates in an $n$-qubit circuit, there are $n$ possible actions, one for each qubit. For the synthesis to be considered exact, the cost function (as specified in Eq. 1) must be exactly zero, only allowing the circuit to induce a global phase difference.
The agent incurs a penalty of $-1$ for every gate added to the circuit and receives a reward of 00 upon exact synthesis of the target unitary. This incentivizes the agent towards shorter episodes and thus minimizes the total number of gates placed [^1] An alternative is to penalize the agent for every $T/T^{\dagger}$ gate added to the circuit. In our experiments, this gave worse total gate count with no improvement in the number of T-gates.. An episode concludes once the target unitary has been successfully synthesized or when a preset maximum number of steps is exceeded. For the observation of the agent, we use $UV_{t}^{\dagger}$ instead of, for example, the concatenation of $U$ and $V_{t}$. This is expected to stabilize training by i) normalizing the goal state observation to the identity matrix independent of the target $U$ and ii) alleviating the curse of dimensionality. Intuitively, $UV_{t}^{\dagger}$ can be interpreted as the unitary residual toward the target unitary $U$, meaning the unitary of the yet unbuilt part of the final circuit.
b) Generation of Target Unitary: To generate target unitaries during training, we initiate each episode by sequentially building a random quantum circuit from the Clifford+T gate set and then calculating the unitary matrix $U$ of the resultant target circuit $C_{U}$. The synthesis task’s complexity is presumed to increase with the increment in the number of gates comprising the target circuit. To facilitate learning, we confine the gate count of sampled circuits to a specified interval with a feasible upper limit.
Calculate:
Target circuit generation
$C_{t}$
$C_{U},U$
$V_{t+1}$
H
T
T
$\text{T}^{\dagger}$
$A_{t}$
$O_{t+1},r_{t+1}$
$M_{t+1}$
H
H
T
T
T
$C_{t+1}$
Gate alphabet
$\text{T}^{\dagger}$
$\text{T}^{\dagger}$
H
• HS cost
• Reward
• Observation
• Action mask
Figure 1: Schematics of the RL environment. An episode begins by sampling a random circuit $C_{U}$ from the Clifford+T gate set to establish the target unitary $U$ for the episode. The state of the episode is initialized with an empty circuit. At each step $t$: (1) The circuit $C_{t}$ is augmented by appending the gate corresponding to action $A_{t}$, yielding a new circuit $C_{t+1}$ with its associated unitary $V_{t+1}$; (2) The target unitary $U$ and the new unitary $V_{t+1}$ are compared using the metric defined in Eq. 1; (3) The reward $r_{t+1}$ is calculated based on the comparison from step (2); (4) Any operations that would lead to gate cancellations or redundancies in $C_{t+1}$ are marked as invalid for the subsequent step; (5) The agent receives the next observation $O_{t+1}$, the reward $r_{t+1}$, and the action mask $M_{t+1}$.
Figure 2: Example of action constraints: The agent has to check the circuit $C_{t-1}$, for various cancellations and redundancies before appending the next gate. A $T^{\dagger}$ gate cannot be place on the first qubit since it commutes all the way to the first moment and cancels the $T$ gate. An $H$ gate cannot be placed on the second qubit since it cancels the already present $H$ gate. This is done until no more cancellation/commutations are possible.
c) Constraining the Search Space: Our approach employs tree search to tackle the challenge of synthesizing unitary operations. In particular, we implement Gumbel AlphaZero to conduct a parallel search, aiming to identify the optimal circuit within the set of potential solutions for a given unitary operation. To alleviate the search complexity, we enforce state-dependent constraints on the set of actions that the agent can select at each time step [41]. More specifically, given that our method is based on a tree search strategy, our goal is to avoid cycles in the state evolution that can arise from the effects of gate cancellations. Consider, for instance, the scenario where two successive gates negating each other are applied, such as an H-gate followed by another H-gate on the same qubit. This sequence of operations would return the system to its prior state, rendering the exploration of such branches in the tree futile. Similarly, our goal is to avoid gate sequences that could be implemented by a single gate from the original gate set. For example, applying a T-gate followed by another T-gate on the same qubit is equivalent to applying a single S-gate.
Given that many gates from the Clifford+T set commute, we assess these conditions for all gates within the circuit that commute to the earliest available circuit moment. More formally, for two unitary gates $A,B\in\mathcal{G}$, with $\mathcal{G}$ denoting the Clifford+T gate set, we say that $A$ and $B$ commute iff $AB=BA$. Assuming gate $A$ is part of the circuit $C_{t-1}$ and commutes to the earliest vacant circuit moment, we invalidate the action of appending a gate $B$ to the circuit at time step $t$ when either of the following conditions is satisfied:
  1. 1.
    Cancellation: $BA=I$, where $I$ is the identity matrix.
  2. 2.
    Redundancy: $BA=C$, where $C\in\mathcal{G}$ is a gate within the gate set.
We provide an example of these constraints in Fig. 2. While not comprehensive, our action masking approach is anticipated to significantly reduce the search complexity. To provide challenging targets for the agent during training, we employ the same masking strategy during the sampling of a target unitary $U$.
VI Experiments
In this section, we present the experimental results of our RL agent for synthesizing unitaries and describe our setup in Section VI-A. We evaluate the agent on two types of unitaries: random unitaries in Section VI-B and structured unitaries in Section VI-C. We compare the performance of our agent with the state-of-the-art synthesis algorithms MIN-T-SYNTH [11] and QuantumCircuitOpt  [41].
VI-A Implementation and Setup
Our environment code was fully implemented in JAX, leveraging its JIT (just-in-time) compilation to address the inherent complexities of the training and inference process and enable seamless agent parallelization on a GPU. To ensure correctness, the calculation of the unitary using our framework for a given circuit were cross-validated with the comprehensive quantum computing software Qiskit [60]. For Gumbel AlphaZero, we utilize an implementation from the DeepMind mctx library [21] for all of our experiments, which is also implemented in JAX.
We employ simple feed-forward neural networks for the actor and critic model. Each network is composed of five layers, with 1024 hidden units per layer, and incorporates layer normalization [61] for regularization purposes [62]. To improve stability during the training of the critic network, we normalized the critic outputs into the range of $[-1,1]$. To process the complex-valued observations $UV_{t}^{\dagger}$, we simply split the real and imaginary parts into separate entries and flatten the resulting matrix.
a) Training: Our agent’s training is conducted over two million time steps, utilizing a replay buffer with the capacity to store 100,000 transitions. Here, we distribute the data collection across 256 concurrent environment workers and compute the training step gradients using a batch size of 1024. As a result, each transition sample has an update-to-data ratio of four. We sample the gate count of target circuits uniformly from the range $[3,\beta]$. The upper limit $\beta$ is set to $60,60,60,40$ for 2,3,4,5 qubit models, respectively. We reduce $\beta$ as the number of qubits increases to alleviate the exponential growth in the size of the observation and action spaces. The gate set of our experiments is $\mathcal{G}=\{H,T,T^{\dagger},S,S^{\dagger},Z,CX\}$.
b) Parallel search using Gumbel AlphaZero: Given a Gumbel AlphaZero model that has been adequately trained, meaning weights $\theta,\mu$ of the policy network $\pi_{\theta}$ and the value network $\upsilon_{\mu}$ are available, we implement a parallelized search strategy. In contrast to sampling actions greedily with Gumbel samples set to zero leveraging a single agent, we follow the same strategy of Gumbel-based action sampling of training with multiple stochastic agents deployed in parallel. From the candidate solutions, we select the one with the highest score, which is determined by a scoring function that evaluates the quality of the solution (either the total gate count or the number of T-gates).
This procedure takes advantage of the variability stemming from the Gumbel distribution to investigate a wider array of possible solutions, increasing the diversity of the search trajectories and improving the likelihood of discovering superior circuits. Further, parallelizing the search not only reduces the chance of sub-optimal solutions but also expedites the search by making efficient use of available computational resources. Most importantly, it allows the natural incorporation of two-level optimization goals within the search process: minimizing the overall gate count, encouraged by a $-1$ reward penalty for every gate added, and minimizing the number of T-gates by scanning the collection of identified solution circuits in a subsequent step. The specifics of our algorithm are presented in Algorithm 1, with the internal data structure of our implementation below it.
Algorithm 1 Parallel search using Gumbel AlphaZero
0:  $n_{a}$: number of actions; $n_{sim}$: number of simulations; $k$: number of actions to sample; $b$: number of runs
0:  $\pi_{\theta}$: policy network; $\upsilon_{\mu}$: value network
0:  $score(x)$: Solution scoring function
0:  $s_{0}$: Initial state of the environment to solve
1:  Replicate initial state $s_{0}$ $b$ times: $\vec{s}\leftarrow[s_{0}^{0},s_{0}^{1},...,s_{0}^{b-1}]$
2:  while not all finished for every non-terminal run $i$ do
3:     $s_{0}\leftarrow\vec{s}_{i}$
4:     Compute logits using $\pi_{\theta}(s_{0})$
5:     Sample Gumbel variables $(g\in\mathbb{R}^{n_{a}})\sim\text{Gumbel}(0)$
6:     Find $k$ actions using the Gumbel-top-$k$ trick
$\mathcal{A}_{\text{topk}}=\mathrm{argtop}\;(g+\text{logits},k)$
7:     Use sequential halving with $n_{sim}$ MCTS simulations, root action selection $g(a)+\text{logits}(a)+\sigma(Q(s_{0},a))$, deterministic non-root action selection and $\upsilon_{\mu}$ for leaf node evaluation
8:     $A_{n_{sim}+1}=\underset{a\in\text{Remaining}}{\mathrm{argmax}}(g(a)+\text{logits}(a)+\sigma(Q(s_{0},a)))$
9:     Step the environment with action $A_{n_{sim}+1}$ and update to successor state $s^{\prime}$: $\vec{s}_{i}\leftarrow s^{\prime}$
10:  end while
11:  Identify runs with valid solutions and their indices $I_{valid}$
12:  return  $\underset{i\in I_{valid}}{\mathrm{argmax}}\;score(\vec{s}_{i})$
Environment Data Structures
1:  Data Structure QuantumCircuit
2:   $circuit\;layout$: List[List[List[integer]]]
3:   $unitary$: List[List[complex]]
4:   $operations\;forward\;commuting$: List[boolean]
5:   $gate\;counts$: List[integer]
1:  Data Structure EnvironmentState
2:   $time\;step$: integer
3:   $circuit$: QuantumCircuit
4:   $action\;history$: List[integer]
5:   $target\;circuit$: QuantumCircuit
6:   $target\;action\;history$: List[integer]
7:   $fidelity$: float
8:   $rng\;key$: integer
Figure 3: Schematics of the Evaluation: The schematic representation of the evaluation of a single circuit performed by our RL agent. We evaluate a given unitary in parallel and choose the resulting circuit with minimal # T-gates/total gates.
c) Evaluation procedure: As noted in the previous paragraph, we leverage parallel search during evaluation. Here, a single target unitary is repeated 256 times, from which the best solution circuit is reported, in our case, circuits with the lowest gate count or the fewest T-gates (depending on the experiment). The schematic for evaluation is described in Fig. 3
VI-B Results for Random Unitaries
For our evaluation of randomized unitaries, we sample 50 target unitaries from randomly generated circuits per data point. We do this using rejection sampling, where we sample random circuits until we find 50 distinct target unitaries with a specific number of T-gates or total number of gates. Specifically, we sample circuits with 0 to 20 T-gates and 3 to 60 total gates for 2,3 and 4 qubits. For 5 qubits, we use the same range of T-gates but limit ourselves to 0 to 40 total gates. For comparison, we evaluate the same circuits using PyZX [63], an optimization heuristic based on ZX-calculus [64], which can reduce the number of T-gates in a circuit. The results of PyZX optimization can be seen as heuristic upper bound on the T-count of a unitary if the corresponding circuit is known. We also compare the total number of gates after PyZX optimisation and our agent’s synthesis.
(a) Comparison of average number of T-gates in the synthesised circuit by our RL agent and the optimised circuit by PyZX.
(b) Distribution of unitaries our RL agent successfully ($C_{HS}<0.1$) synthesised out of the 50 unitaries.
Figure 4: Our RL agent’s performance on synthesizing 50 unitaries, generated by randomly sampling Clifford+T gate set, building a circuit and calculating the corresponding unitary. Here, the number of T-gates for each unitary is fixed and ranges from 0 to 20 for 2, 3 and 4 qubits and 0 to 15 for 5 qubits. (a) Our agent can synthesize 2 and 3 qubit unitaries with similar average T-gate count as PyZX. PyZX performs slightly better on average for 4 and 5 qubit unitaries. (b) We see a slight decline in performance for higher T-gate count for 2 and 3 qubit unitaries, and a significant drop for 4 and 5 qubit unitaries. The observed decline in performance can be attributed to the imbalance in the training set distribution. Specifically, there are fewer instances of circuits with larger T-gate counts compared to those with smaller counts.
In Fig. 4 we plot the performance of our RL agent in synthesizing random unitaries for a fixed number of T-gates. Fig. 4(a) shows the average and median number of T-gates obtained by our RL agent over 50 random unitaries. Here, the line $x=y$ denotes the case where the number of T-gates in the input and the output circuits are the same. In Fig. 4(b), we plot the fraction of unitaries the agent successfully synthesized. The success is determined by calculating $C_{HS}$ (Eq. 1), for each of the 50 instances. When $C_{HS}>0.1$, evaluation is considered failed and $C_{HS}<0.1$ is considered success [^2] Since PyZX is not a synthesis tool there is no possibility of failure for such small qubit count. PyZX is always able to either reduce the number of T-gates in the circuit or at least keep it the same.. We empirically found the threshold of $0.1$ to work best for our experiments. Consequently, we verified whether the results with $C_{HS}<0.1$ match the target unitary, which we found to be the case in all the experiments. The error band shown in Fig. 4(a) around the RL agent curve is the fraction of unsuccessful runs out of the 50 unitaries considered. We see that the error band broadens for a higher number of T-gates, showing a decrease in performance. In Fig. 5, we show the performance for synthesizing unitaries for a fixed number of total gates.
(a) Comparison of the average total gate count in the synthesized circuit by our RL agent and the optimized circuit by PyZX.
(b) Distribution of unitaries our RL agent successfully ($C_{HS}<0.1$) synthesised out of the 50 unitaries.
Figure 5: Our RL agent’s performance on synthesizing 50 unitaries, generated by randomly sampling Clifford+T gate set, building a circuit and calculating the corresponding unitary. Here, the total number of gates for each unitary is fixed and ranges from 3 to 60 for 2, 3 and 4 qubits and 3 to 40 for 5 qubits. (a) Our agent can synthesize 2,3 and 4 qubit unitaries with better average total gate count than PyZX, although we see a drop in performance for higher total gate count. Average total number of gates is similar for the case of 5 qubit unitaries. (b) We see a slight decline in performance for 2 and 3 qubit unitaries, and a significant drop for 4 and 5 qubit unitaries for larger gate count.
We find that our RL agent can synthesize all the target unitaries for small T-gate count and total gate count. The decrease in performance for T-gate count, can be attributed to the skewed distribution of T-gate numbers in the training set, meaning there are fewer circuits with larger T-gate counts given the current upper limit of 60 total gates. The performance decline for larger number of qubits can be attributed to the growing number of possibilities for gate placement on the qubit lines along with higher dimensional observation vectors (number of elements of the unitary matrix). For 3,4, and 5-qubit unitaries, our RL agent performs similarly to PyZX in terms of the number of T-gates in the optimized circuit. The reduction in total gate count using our RL agent was much better than PyZX since it has been observed [65] that PyZX T-gate optimization can increase the total number of gates. This highlights the benefits of our two-level optimization (both on the number of T-gates and total gates) procedure, which targets both the number of T-gates and the total gate count.
VI-C Evaluation on Structured Unitaries
Table I: Benchmarking our method on structured unitaries against MIN-T-SYNTH [11] and QCOpt  [41]. We compare the number of T-gates and the run-time of synthesis. The success ratio (as defined in Fig. 3 and Section VI-B (b)), in the column for our RL agent, provides the number of successful runs (‘-’ indicates algorithm was unsuccessful in finding the synthesis either due to time limit or no convergence to the solution). The example decompositions of some unitaries are given in [10].
Qubits Unitary Our RL agent MIN-T-SYNTH QCOpt
T Time(s) T Time(s) T Time(s)
2 CP 3 10.69 3 0.007 3 6.15
CV 3 9.82 - - 3 4.16
CY 0 9.50 0 0.224 3 1.89
CZ 0 9.41 0 0.330 0 0.1
SWAP 0 9.27 0 0.353 0 0.01
W 2 $\bm{9.62}$ 0 3221.21
CH 2 $\bm{16.23}$ 2 147.12
CT - - - - - -
3 Toffoli 7 25.92 7 4.99 7 10800
Single
Negated
Toffoli
7 16.83 7 5.70 - -
Double
Negated
Toffoli
7 16.78 7 3.39 - -
Fredkin 7 16.96 7 4.85 - -
Peres 7 16.87 7 5.00 - -
QOR 7 16.56 7 5.09 - -
AND 7 16.83 7 3.44 - -
Miller 7 16.61 7 3.47 - -
TR 7 16.19 7 3.47 - -
4
Toffoli
(1 ancilla)
7 $\bm{31.50}$ 7 233.54 - -
2 Toffoli
12 $\bm{19.27}$ 11 6642.16 - -
3 Toffoli
7 $\bm{18.54}$ 7 309.84 - -
1-bit adder 7 $\bm{18.53}$ 7 312.55 - -
2 Peres 7 $\bm{18.85}$ 7 312.55 - -
For this experiment, we selected unitaries that are commonly encountered in various algorithms. These unitaries, such as Toffoli, Fredkin, controlled-Phase (CP), and controlled-$\sqrt{X}$ (CV) gates, are crucial for decomposing multi-qubit multi-controlled unitaries efficiently [66].
We compare our agent’s performance with other synthesis algorithms and depict the results in Table I. The table includes the number of qubits in the input circuit, the target unitary, the results obtained by our agent, and the results of MIN-T-SYNTH and QuantumCircuitOpt (QCOpt for short). For MIN-T-SYNTH, we used the code provided by the authors of Ref. [11]. While the evaluation was successful for some of our chosen examples, manual code adjustments would have been required, leading to ‘-’ entries in the table. For QCOpt, the method exceeded the time limit of three hours without finding a solution in the basic Clifford+T gate set, which are also denoted by ‘-’ in the table.
Our method successfully synthesized all the unitaries, except for the controlled-T gate, which was not synthesized by any of the algorithms. To the best of our knowledge, it is unknown whether an exact synthesis is even possible, without the use of ancilla qubits as per Theorem 2 in Ref. [39]. As per Refs. [10, 11], we our RL agent achieves optimal of T-count for all target unitaries except $U_{1}$, which was synthesized with one extra T-gate compared to MIN-T-SYNTH.
Both QCOpt and our RL agent require an explicitly defined input gate set, which for $n$-qubit unitary would require $\mathcal{O}(n^{2})$ possible ways to place CNOT gates and $\mathcal{O}(n)$ one-qubit gates unless we already have prior knowledge about the decomposition of the unitary. The difference between QCOpt and our method is that the mixed-integer programming (MIP) method that QCOpt uses becomes highly inefficient in terms of run-time under such scaling. In comparison, our agent only has to account for these possibilities during training.
Comparing the run-times of our agent with the heuristic methods, we observed that, while both MIN-T-SYNTH and QCOpt are faster for simple unitaries, our agent takes approximately fifteen seconds for synthesis. However, as we analyze the increase in run-time with the number of qubits, we notice a significant increase in synthesis time for the other methods, whereas our method remains relatively constant. Therefore, the most advantageous aspect of our RL algorithm is the inference time as we scale to higher qubits.
Our results are promising as they did not require handcrafted rules to handle the complexities of search-tree pruning and gate placements. Instead, we only leverage commutation rules and gate cancellation identities. Throughout training, our agent learns to account for these rules under the single incentive to reduce the total gate count.
VII Outlook
In our presentation of our RL framework, we focused on the exact synthesis of unitaries using the Clifford+T gate set. With some further improvements, the scope of our agent’s application can be expanded. Firstly, our approach does not preserve the MCTS tree following the execution of actions, meaning the search tree is rebuilt from scratch at every iteration. We aim to address this limitation in our future work. Moreover, akin to QCOpt, augmenting the action masking process to encompass additional circuit equivalence relations might be feasible. A further enhancement could involve adopting a more advanced neural network design, like a transformer network, which could consider the circuit’s configuration in addition to the target unitary, potentially elevating the agent’s efficacy. An intriguing inquiry is whether a more effective method exists for representing the unitary as input to the neural network.
Transitioning to approximate synthesis theoretically entails dropping the constraint that $C_{HS}$ be precisely zero for two unitaries. Another consideration is to incorporate unitaries that utilize ancilla qubits. This is because certain unitaries, like the controlled-T, may necessitate an ancilla qubit for their synthesis. Additionally, using an ancilla qubit could reduce the depth of the synthesized circuit.
VIII Conclusion
Traditional heuristic approaches to unitary synthesis rely on manually designed rules to reduce search space effectively. On the other hand, automated methods utilizing RL endure the initial cost of extensive training periods but can independently discover these synthesis strategies. Our empirical findings indicate that the inference times did not markedly increase with the number of qubits, marking a notable improvement over traditional heuristic techniques. Therefore, our proposed method establishes a new benchmark for unitary synthesis algorithms that employ reinforcement learning.
Acknowledgement
We are very grateful to P.  Mukhopadhyay for helpful discussions and for sharing the code with us, which was used to generate the results for the MIN-T-SYNTH method shown in Table I. We would also like to thank V. Gheorghiu for sharing the Github repository with us for the approximate circuit synthesis in Clifford+T gates. We thank N. Quetschlich for interesting discussions related to quantum compilation using reinforcement learning. Finally, we would like to thank Q. Göttl for comments on the manuscript.